Please be thorough so I can learn it! If you're correct, I'll rate you lifesaver!
Thanks!!

A point charge of +6.8µ C is located 0.29 m from a second point charge of -4.9µ C. What is the absolute magnitude of the force on each charge?
Answer #1

A point charge of +6.8µ C is located 0.29 m from a second point charge of -4.9µ C. What is the absolute magnitude of the force on each charge?
F=?
Q1=+6.8 µ C
Q2=-4.9 µ C
D=0.29m
K=9*10^9 N m^2/C^2
According to Coulomb’s law
F=K*q1*q2/d^2
F=(9*10^9*6.8*10^-6*(-4.9)*10^-6)/(0.29^2)=3.565 N
Answer #2

According to Coulomb's law: